a crowbar is 4.5 feet long. where must the fulcrum be placed so that a weight of 1600 pounds can be moved by exerting a force of 200 pounds?
Answer by checkley77(12844)

(4.5-x)/x=200/1,600
200x=1,600(4.5-x)
200x=7,200-1,600x
200x+1,600x=7,200
1,800x=7,200
x=7,200/1,800
x=4 ft. is the distance from the 200 pound force.
4.5-4=.5 ft. from the 1,600 pound load.
Proof:
4*200=.5*1,600
800=800